Listed below are the numbers of words spoken in a day by each member of eight different randomly selected couples. Complete parts​ (a) and​ (b) below.

Male

16,054

27,338

1411

8019

19,507

15,326

13,597

25,343

Female

25,186

13,173

18,067

17,119

13,192

17,360

16,313

18,554

a. Use a

0.05

significance level to test the claim that among​ couples, males speak fewer words in a day than females.In this​ example,

μd

is the mean value of the differences d for the population of all pairs of​ data, where each individual difference d is defined as the words spoken by the male minus words spoken by the female. What are the null and alternative hypotheses for the hypothesis​ test?

H0​:

μd

equals=

00 ​word(s)

H1​:

μd

less than<

00 ​word(s)

​(Type integers or decimals. Do not​ round.)

Identify the test statistic.

t=nothing

​(Round to two decimal places as​ needed.)

Identify the​ P-value.

​P-value=nothing

​(Round to three decimal places as​ needed.)

What is the conclusion based on the hypothesis​ test?

Since the​ P-value is

greater than

less than or equal to

the significance​ level,

fail to reject

reject

the null hypothesis. There

is not

is

sufficient evidence to support the claim that males speak fewer words in a day than females.

b. Construct the confidence interval that could be used for the hypothesis test described in part​ (a). What feature of the confidence interval leads to the same conclusion reached in part​ (a)?

The confidence interval is

nothing

​word(s)<μd<nothing

​word(s).

​(Round to the nearest integer as​ needed.)

What feature of the confidence interval leads to the same conclusion reached in part​ (a)?

Since the confidence interval contains

zero,

only negative numbers,

only positive numbers,

reject

fail to reject

the null hypothesis.
Answer #1

Assuming that the data given in the question is as follows :

Male Female Difference (D)
16,054 25,186 -9,132
27,338 13,173 14,165
1411 18,067 -16,656
8019 17,119 -9,100
19,507 13,192 6,315
15,326 17,360 -2,034
13,597 16,313 -2,716
25,343 18,554 6,789

Thus,  \bar{x}_{D} = -1546.125, SD = 10160.37487, n = 8

Ho :  \mu_{D} = 0
Ho :  \mu_{D} < 0

Level of Significance(l.o.s.) :  \alpha = 0.05

Decision Criteria : Reject Ho at 5% l.o.s. if t cal < t tab,
where t tab = t (\alpha, n-1) = t (0.05,7) = - 1.8946

Calculation :

t cal =  \frac{\bar{x}_{D}-\mu_{D}}{S_{D}/\sqrt{n}} =  \frac{-1546.125-0}{10160.37487/\sqrt{8}} = -0.43041 = -0.43

p-value = P( t < -0.43 ) = 0.340061 = 0.340

Conclusion : Since the​ P-value is less than the significance​ level, we fail to reject the null hypothesis. There is not
sufficient evidence to support the claim that males speak fewer words in a day than females.

b) A (1-\alpha)% confidence interval for population mean difference is obtained as follows :
(\bar{x}_{D}\mp t_{\alpha /2,n-1}*S_{D}*\sqrt{\frac{1}{n}})
Thus, 95% confidence interval for population mean difference is :
(-1546.125-1.8946*10160.37487*\sqrt{\frac{1}{8}},-1546.125+1.8946*10160.37487*\sqrt{\frac{1}{8}})

\Rightarrow ( -8351.9734, 5259.7234 )
\Rightarrow ( -8352, 5260 )

Since the confidence interval contains zero we fail to reject the null hypothesis.

Hope this answers your query!